OpenStudy (anonymous):

find the values for the trigonometric functions of t from the given information. tant= -3/4, cost >0

OpenStudy (lgbasallote):

i'll let my assistant @waterineyes take this one :)

OpenStudy (anonymous):

Do you want to find cost sint cott etc etc all the values @tamy

OpenStudy (anonymous):

yes, all those values. thanks for helping =]

OpenStudy (anonymous):

So we have to use some identities here; \[\large \sin^2 \theta + \cos^2 \theta = 1\] \[\large \cot \theta = \frac{1}{\tan \theta}\] \[\large 1 + \tan^2 \theta = \sec^2 \theta\]

OpenStudy (anonymous):

I think they will be sufficient for the solution..

OpenStudy (lgbasallote):

my assistant is very good :DDD haha lol

OpenStudy (anonymous):

Now let us find sec(t) from the third identity I gave you.. \[\large \sec^2(t) = 1 + (\frac{-3}{4})^2 \implies 1 + \frac{9}{16} \implies \frac{25}{16}\] So now take the square root: \[\large \sec(t) = \frac{5}{4}\] Getting till here @tamy

OpenStudy (anonymous):

And you know: \[\large \cos(t) = \frac{1}{\sec(t)} \implies \frac{1}{\frac{5}{4}} \implies \cos(t) = \frac{4}{5}\]

OpenStudy (anonymous):

Now using first identity I gave you: \[\large \sin^2 (t) = 1- \cos^2 (t) = 1 - (\frac{4}{5})^2 = 1 - \frac{16}{25} \implies \frac{9}{25}\] Now take the square root: \[\large \sin(t) = \frac{3}{5}\]

OpenStudy (anonymous):

Can you do further or not ?? @tamy

OpenStudy (anonymous):

yes I think I got it. Thank u so much!

OpenStudy (anonymous):

For your easiness I explain you more: \[\large cosec(t) = \frac{1}{\sin(t)}\] By using this find cosec(t).. \[\large \cot(t) = \frac{1}{\tan(t)}\] By using this find cot(t)..

OpenStudy (anonymous):

@tamy try it once and if you have problem then let me know..

OpenStudy (anonymous):

I will if I do. u are very helpful thanks =]

OpenStudy (anonymous):

Welcome dear..